A horizontal square wire loop of side d lies in a plane perpendicular to a uniform magnetic field B pointing from above into the plane of the loop. The magnitude of magnetic eld ~ B is: B = 0.80 T. If in ∆t = 4.0 ms the wire is reshaped from a square to a rectangle of dimensions: L = 10.0 cm; W = 6.0 cm, but remains in the same plane. a - What is the side d of the square? b - What is the magnitude of the average induced emf, in the wire during this time?
Answer #1

Ba0.80T Af=y.oxioss La loom w=6cm @ we know that in this questim l perimeter of square)= (perimeter of rectas) 4(a)=2(L+w) 49② we know that E = - doo [(-) sign show that entdian is antiof magnetic fields] d[ B.A E Es B ( Arenog(szundred E = B(a? (w)E B ((9) Cloth )) E B (4* af Es 0.80(64-60) ToxuX10-3 E=0.8074 Чл|- | E- 0.8x10l volt E = 8 volt cs Scanne - Camscanner